OpenStudy (anonymous):

The ratio of men to women in a certain factory is 3 to 4. There are 222 men. How many workers are there?

OpenStudy (freckles):

\[\frac{m}{w}=\frac{3}{4} \\ m=222\]

OpenStudy (freckles):

solve for w

OpenStudy (freckles):

then find w+m

OpenStudy (freckles):

m to w is 3 to 4 is how I setup the first equation m is 222 is how i got the second equation

OpenStudy (eric_d):

4/3 x 222 + m

OpenStudy (eric_d):

= w

OpenStudy (eric_d):

Get me @jdmeduardo ?

OpenStudy (eric_d):

there are 222 men

OpenStudy (eric_d):

women=(4/3)*222

OpenStudy (eric_d):

Add them up to get the total number of workers @jdmeduardo

OpenStudy (anonymous):

yes thank you

OpenStudy (eric_d):

You're welcome @jdmeduardo